OpenStudy (anonymous):

Please Help! If f(x) = x-2/2 , then what is f^–1(x) ? I am trying to help my little brother on this question, it has been so long that I have forgotten how to do these type of problems.

OpenStudy (anonymous):

These are the answers... a. 2x + 2 b. 2x – 1 c. 2x – 2 d. 1/2x-1

OpenStudy (potatoes.ramu):

That is 1/(f(x))

OpenStudy (potatoes.ramu):

So you'll get the reciprocal of f(x)

OpenStudy (anonymous):

no actually it is the inverse function, not the reciprocal

OpenStudy (anonymous):

@Potatoes.ramu Do you mind if you explain a more? I'm not really catching on.

OpenStudy (anonymous):

put \[x=\frac{y-2}{2}\] and solve for \(y\)

OpenStudy (zale101):

f(x) = (x - 2) / 2 let f(x) be y y = (x - 2) / 2 since you're looking for f^-1, it means to look for the inverse, which means to interchange y to x and x to y x = (y - 2) / 2 multiply both sides by 2 2x = y - 2 add 2 to both sides 2x + 2 = y now switch the y for f^-1(x) f^-1(x)=2x+2

OpenStudy (anonymous):

@bobers go with @Zale101 answer, it is the correct one

OpenStudy (anonymous):

@satellite73 Thank you.